# Absorption Spectroscopy of 40Ca Atomic Beams Produced via Pulsed Laser Ablation: A Quantitative Comparison of Ca and CaTiO3 Targets

### v1

@@ Abstract
Pulsed laser ablation is an increasingly prevalent method for fast ion trap loading of various species, however characteristics of the ablation target source material can affect the ion-loading process. One factor which can reduce the atomic flux from a target is oxidation during atmospheric exposure when preparing or making changes to the ion trap vacuum system. Recent work has shown that perovskite ablation targets produce consistent atomic densities even after exposure to atmosphere when compared to elemental source targets. In this work, we directly compare calcium (Ca) and calcium-titanate (CaTiO3) ablation targets, characterizing the neutral atomic beam flux using resonant, time-resolved absorption spectroscopy of the 423 nm transition in neutral Ca. We measure the ablation plume longitudinal and transverse temperatures, number density, ion production, and spot lifetime for each target. In addition, we compare the ablated atomic beam density for both targets before and after 21 hours of exposure to atmosphere, demonstrating the relative robustness of the CaTiO3 source.
@@ 1 Introduction
Increasing numbers of researchers are moving towards laser ablation loading of trapped ions for many practical reasons including fast loading times [14, 11] and elimination of residual trap heating. Several groups have demonstrated successful ablation loading of various species of ions in RF surface and Penning ion traps [14, 13, 12, 17, 11, 18, 19, 20, 21]. A primary concern in laser ablation loading of ions is the atom yield per ablation pulse and the number density stability versus ablation pulse number (i.e. spot lifetime) and thus how often the ablation laser spot must be re-positioned on the target. As with oven sources, oxidation of the target sample will reduce the available atom number density (e.g. during prolonged trap assembly or subsequent vacuum system breaks). Several groups have shown ablation loading targets to have relatively long spot lifetimes of tens of thousands of pulses given laser fluences in the 0.1 - 3.0 J/cm2 range[14, 13, 12, 22, 19, 21]. Additionally, laser ablation targets of different salt and perovskite substrates have been shown to produce a consistent ion yield when compared to elemental sources which are subject to oxidation when exposed to atmosphere [23, 24, 22].
In this work, we characterize 40Ca neutral atom plumes produced from pulsed laser ablation of both Ca and CaTiO3 targets for future integration with, for example, a compact Penning ion trap [25, 26, 27]. Calcium metal is highly reactive with oxygen and an oxidative layer is formed when the metal is exposed to atmosphere. However, CaTiO3 is chemically inert and does not degrade under air exposure, making it a potential alternative as a 40Ca ion source in ion-trapping experiments. In this work, we perform a comparative characterization of the neutral atomic flux produced through pulsed laser ablation of Ca and CaTiO3 targets using absorption spectroscopy of the 423 nm transition in neutral Ca. We derive an expression for the neutral atomic density extracted from resonant laser beam absorption measurements, giving quantitative comparisons of the ablation plume temperature and spot lifetimes under vacuum as well as the ground state atomic densities for both targets before and after exposure to atmosphere. We also measure the time-dependent phase space profiles of ablation plumes in the absence of restrictive trap electrode geometries. Combined with photoionization or electron impact ionization cross sections, our temperature and density measurements permit quantitative estimates of 40Ca ablation trap loading rates for arbitrary trap geometries.
@@ 3.1 Ablation Plume Temperature and Density Analysis
To extract the longitudinal plume temperature, we fit the plume velocity distributions for a range of ablation laser fluences to a Boltzmann distribution, which can be seen in Fig. 3. We observe most probable velocities for the range of fluences shown in Fig. 3 of 1070-1425 m/s and 1332-1562 m/s for the Ca and CaTiO3 targets respectively, which agrees well with previous pulsed laser ablation experiments [11, 13, 19]. The velocity profiles follow Boltzmann distributions with the exception of the highest-velocity atoms in the high-pulse-fluence regime. This non-thermal atomic beam component may be the result of collisional effects at higher ablation laser fluence, which have been shown to cause accelerated expansion of the metal vapor from the target surface [31]. In Fig. 4(a), we compare the extracted longitudinal temperature vs. ablation laser fluence for each target, which exhibits a qualitatively linear dependence. Similarly, we plot the calculated peak atomic number density vs. laser fluence for each target in Fig. 4(b).
@@ 3.2 Spot Lifetime
One concern with the use of ablation targets for loading of trapped ions is the ablation spot lifetime. During the early stages of our experiment, we observed depletion of ablation spots at high fluences while initially setting up the absorption spectroscopy. To quantify the spot lifetime as a function of laser fluence, we place the probe laser beam at the peak plume density vertical position () near the target surface and increase the ablation laser beam fluence to nearly extinguish of the probe laser beam transmission signal. We also increase the repetition rate of the ablation laser to the maximum 15 Hz. We collect measurements of the absorption signal over a timescale of approximately 20 minutes. Converting the absorption signal to peak atomic number density, we quantify the peak atomic density as a function of ablation pulse number. As seen in Fig. 4(d), we plot the density vs. number of ablation pulses for each target and fit these to a single-exponential function to extract decay constants for the spot lifetimes. We measure similar 1/ decay pulse numbers of and for the Ca and CaTiO3 targets, respectively. By contrast, Osada et al. recently reported CaTiO3 exhibiting a longer ablation target lifetime when compared to a pure Ca target for a 1064 nm ablation laser wavelength [24]. This could be the result of various differences in our experimental parameters including ablation laser wavelength, angle of incidence, or fluence [32, 33, 34, 35, 36].
@@ 3.3 Target Oxidation
An additional concern that arises in trapped ion loading sources is oxidation of the source when making vacuum system changes. Atmospheric exposure typically causes oxidation of elemental alkaline earth sources. To test the resistance of CaTiO3 to oxidation, we directly compare the plume densities from Ca and CaTiO3 targets under similar laser ablation fluences both before and after exposure to atmosphere. We first measure the velocity distributions for each target under ideal conditions in-vacuum as discussed earlier in Sec. 3. After obtaining the nominal velocity distributions, we then break vacuum and allow the chamber to sit at atmosphere for approximately 21 hours before pumping back down to similar vacuum pressures and repeating the measurement. Figure 5 shows the result of these measurements for both targets. In Fig. 1(a), it is evident that a layer of white oxidation has built up on the Ca target surface after atmospheric exposure, while no visible change was observed for the CaTiO3 target. We observe no reduction in the atomic number density yield of the CaTiO3 target after the atmospheric exposure, however we see an approximately 90% decrease in the yield of the Ca target. These results confirm that CaTiO3 is a good candidate for 40Ca+ experiments requiring frequent vacuum changes, or in complex experimental apparatuses where the vacuum system cannot be immediately pumped after oven assembly (e.g. some cryogenic systems).
@@ 4 Conclusions
We have characterized the phase space distribution of 40Ca neutral atomic beams produced by pulsed laser ablation of Ca and CaTiO3 targets through absorption spectroscopy. Measuring time-dependent absorption of a 423 nm probe laser beam, we observe peak atomic densities from each ablation target on the order of cm-3 at ablation laser fluences in the range of 0.1-0.3 J/cm2. We quantify the ablation plume temperature and velocity for a range of ablation laser pulse fluences, with longitudinal plume temperatures of 2402-4789 K for Ca and 3836-5631 K for CaTiO3. The measured longitudinal plume velocities for the same range of fluences are 1070-1425 m/s and 1332-1562 m/s for the Ca and CaTiO3 targets, respectively. Transverse plume temperatures extracted from measurements of the Doppler broadening of the Ca absorption are in good agreement with the longitudinal temperatures.
Our results highlight the benefit of choosing a perovskite ablation target for ion loading by directly comparing the effects of atmospheric exposure on atomic density yield for each target. We measured minimal decrease in the neutral atomic flux from the CaTiO3 target after 21 hours of atmospheric exposure, while we observed almost complete depletion of the atomic flux from the Ca target after visible surface oxidation of the target. In addition, we quantified the spot lifetime for each target at reasonably high ablation laser fluences (0.18-0.30 J/cm2) where we saw 20-30% decrease in neutral atomic flux after tens of thousands of ablation pulses. We anticipate that the ablated atom plume densities and temperatures reported here, when combined with ionization models, may be used for quantitative estimates of ion trap loading rates in more restricted trap geometries. Future work could include measurements of ablated Ca density within an ion trap apparatus and the relative photoionization loading efficiency of Ca+ obtained from different target materials.

### v2

@@ Abstract
Pulsed laser ablation is an increasingly prevalent method for fast ion trap loading of various species, however characteristics of the ablation target source material can affect the ion-loading process. One factor which can reduce the atomic flux from a target is oxidation during atmospheric exposure when preparing or making changes to the ion trap vacuum system. Recent work has shown that perovskite ablation targets produce consistent atomic densities even after exposure to atmosphere when compared to elemental source targets. In this work, we directly compare calcium (Ca) and calcium-titanate (CaTiO3) ablation targets, characterizing the neutral atomic beam flux using resonant, time-resolved absorption spectroscopy of the 423 nm transition in neutral Ca. We measure the ablation plume longitudinal and transverse temperatures, number density, ion production, and spot lifetime for each target. In addition, we compare the probe laser beam absorption for both targets before and after 21 hours of exposure to atmosphere, demonstrating the relative robustness of the CaTiO3 source.
@@ 1 Introduction
Increasing numbers of researchers are moving towards laser ablation loading of trapped ions for many practical reasons including fast loading times [14, 11] and elimination of residual trap heating. Some commonly-used ablation techniques for ion trap loading are photoionization of neutral atoms within the ion trap [14, 13, 17, 11, 18, 19, 20] or direct ion loading [12, 21], which several groups have demonstrated for various species in RF and Penning ion traps. A primary concern in laser ablation loading of ions is the atom yield per ablation pulse and the number density stability versus ablation pulse number (i.e. spot lifetime) and thus how often the ablation laser spot must be re-positioned on the target. As with oven sources, oxidation of the target sample will reduce the available atom number density (e.g. during prolonged trap assembly or subsequent vacuum system breaks). Several groups have shown ablation loading targets to have relatively long spot lifetimes of tens of thousands of pulses given laser fluences in the 0.1 - 3.0 J/cm2 range[14, 13, 12, 22, 19, 20]. Additionally, laser ablation targets of different salt and perovskite substrates have been shown to produce a consistent ion yield when compared to elemental sources which are subject to oxidation when exposed to atmosphere [23, 24, 22].
In this work, we characterize 40Ca neutral atom plumes produced from pulsed laser ablation of both Ca and CaTiO3 targets for future integration with, for example, a compact Penning ion trap [25, 26, 27]. Calcium metal is highly reactive with oxygen and an oxidative layer is formed when the metal is exposed to atmosphere. However, CaTiO3 is chemically inert and does not degrade under air exposure, making it a potential alternative as a 40Ca ion source in ion-trapping experiments. In this work, we perform a comparative characterization of the neutral atomic flux produced through pulsed laser ablation of Ca and CaTiO3 targets using absorption spectroscopy of the 423 nm transition in neutral Ca. We derive an expression for the neutral atomic density extracted from resonant laser beam absorption measurements, giving quantitative comparisons of the ablation plume temperature and spot lifetimes under vacuum as well as the probe laser beam absorption for both targets before and after exposure to atmosphere. We measure the time-dependent phase space profiles of ablation plumes in the absence of restrictive trap electrode geometries. Combined with photoionization or electron impact ionization cross sections, our temperature and density measurements permit quantitative estimates of 40Ca ablation trap loading rates for arbitrary trap geometries.
@@ 3.1 Ablation Plume Temperature and Density Analysis
To extract the longitudinal plume temperature, we fit the plume velocity distributions for a range of ablation laser fluences to a Boltzmann distribution, which can be seen in Fig. 3. We observe most probable velocities for the range of fluences shown in Fig. 3 of 1502-2215 m/s and 1880-2365 m/s for the Ca and CaTiO3 targets respectively, which agrees well with previous pulsed laser ablation experiments [11, 13, 19]. The velocity profiles follow Boltzmann distributions with the exception of the highest-velocity atoms in the high-pulse-fluence regime. This non-thermal atomic beam component may be the result of collisional effects at higher ablation laser fluence, which have been shown to cause accelerated expansion of the metal vapor from the target surface [31]. In Fig. 4(a), we compare the extracted longitudinal temperature vs. ablation laser fluence for each target, which exhibits a qualitatively linear dependence. Similarly, we plot the calculated peak atomic number density vs. laser fluence for each target in Fig. 4(b).
@@ 3.2 Spot Lifetime
One concern with the use of ablation targets for loading of trapped ions is the ablation spot lifetime. During the early stages of our experiment, we observed depletion of ablation spots at high fluences while initially setting up the absorption spectroscopy. To quantify the spot lifetime as a function of laser fluence, we place the probe laser beam at the peak plume density vertical position () near the target surface and increase the ablation laser beam fluence to nearly extinguish the probe laser beam transmission signal. We also increase the repetition rate of the ablation laser to the maximum 15 Hz. We collect measurements of the absorption signal over a timescale of approximately 20 minutes. Converting the absorption signal to peak atomic number density, we quantify the peak atomic density as a function of ablation pulse number. As seen in Fig. 4(d), we plot the density vs. number of ablation pulses for each target and fit these to a single-exponential function to extract decay constants for the spot lifetimes. We measure similar 1/ decay pulse numbers of and for the Ca and CaTiO3 targets, respectively. By contrast, Osada et al. recently reported CaTiO3 exhibiting a longer ablation target lifetime when compared to a pure Ca target for a 1064 nm ablation laser wavelength [24]. This could be the result of various differences in our experimental parameters including ablation laser wavelength, angle of incidence, or fluence [32, 33, 34, 35, 36].
@@ 3.3 Target Oxidation
An additional concern that arises in trapped ion loading sources is oxidation of the source when making vacuum system changes. Atmospheric exposure typically causes oxidation of elemental alkaline earth sources. To test the resistance of CaTiO3 to oxidation, we directly compare probe laser beam absorption for Ca and CaTiO3 targets under similar laser ablation fluences both before and after exposure to atmosphere. We first measure the time-resolved probe beam absorption signal for each target under ideal conditions in-vacuum as discussed earlier in Sec. 2. After obtaining the nominal probe beam absorption, we then break vacuum and allow the chamber to sit at atmosphere for approximately 21 hours before pumping back down to similar vacuum pressures and repeating the measurement. Figure 5 shows the result of these measurements for both targets. In Fig. 1(a), it is evident that a layer of white oxidation has built up on the Ca target surface after atmospheric exposure, while no visible change was observed for the CaTiO3 target. We observe a 50% reduction in probe beam absorption with the CaTiO3 target after the atmospheric exposure. This reduction in probe absorption is consistent with observed differences in the yield of different ablation spots on the CaTiO3 target. By contrast, we see no evidence of atomic absorption when ablating the Ca target after atmospheric exposure. We also do not observe any noticeable increase in ablated neutral Ca density after 500 ablation pulses. These results confirm that CaTiO3 is a good candidate for 40Ca+ experiments requiring frequent vacuum changes, or in complex experimental apparatuses where the vacuum system cannot be immediately pumped after oven assembly (e.g. some cryogenic systems).
@@ 4 Conclusions
We have characterized the phase space distribution of 40Ca neutral atomic beams produced by pulsed laser ablation of Ca and CaTiO3 targets through absorption spectroscopy. Measuring time-dependent absorption of a 423 nm probe laser beam, we observe peak atomic densities from each ablation target on the order of cm-3 at ablation laser fluences in the range of 0.1-0.3 J/cm2. We quantify the ablation plume temperature and velocity for a range of ablation laser pulse fluences, with longitudinal plume temperatures of 5423-11804 K for Ca and 8499-13453 K for CaTiO3. The measured longitudinal plume velocities for the same range of fluences are 1502-2215 m/s and 1880-2365 for the Ca and CaTiO3 targets, respectively. Transverse plume temperature extracted from measurements of the Doppler broadening of the Ca absorption are below the measured longitudinal temperatures, reflecting the directionality of the atom plumes.
Our results highlight the benefit of choosing a perovskite ablation target for ion loading by directly comparing the effects of atmospheric exposure on atomic density yield for each target. We measured far less reduction in the neutral atomic flux from the CaTiO3 target after 21 hours of atmospheric exposure, while we observed no evidence of atomic flux from the Ca target after visible surface oxidation of the target. The continuous 21-hour exposure is representative of the CaTiO3target longevity when subjected to multiple atmospheric exposures during vacuum changes which can be as short as 1-2 hours. In addition, we quantified the spot lifetime for each target at reasonably high ablation laser fluences (0.18-0.30 J/cm2) where we saw 20-30% decrease in neutral atomic flux after tens of thousands of ablation pulses. We anticipate that the ablated atom plume densities and temperatures reported here, when combined with ionization models, may be used for quantitative estimates of ion trap loading rates in more restricted trap geometries. Future work could include measurements of ablated Ca density within an ion trap apparatus and the relative photoionization loading efficiency of Ca+ obtained from different target materials.
@@
Declarations
Data will be made available upon reasonable request.